# MGAS: Multi-Granularity Architecture Search for Trade-Off Between Model Effectiveness and Efficiency

### v3

@@ IV-C Multi-Stage Search
Let the the search space of the whole architecture be N. To realize the multi-stage search, we decompose $N$ into sequential n moderate-size search spaces $N_{1},...,N_{i},...,N_{n}$ along the network depth, where $N_{i+1}$ is originally connected with $N_{i}$. In this way, the super-net constructed from the search space is naturally separated into n sub-nets with much smaller sizes. We first coarsely search for the sub-net structures in various stages, removing most of the unimportant units, and then fine-tune the remaining architecture until the target model size is achieved.
@@
Input: dataset $D$, target model size $Param_{min}$, sub-net number $n$, target sub-net size $Param_{min}^{s}$ ($s=0,1,...,n$), pruning interval $e$, regrowing ratio $r$, units in the search space $U$
@@
To supervise the optimization of the sub-net, we add task-specific output layers with predefined structure on top of each sub-net to generate results, so that the ground truth can be directly used as the supervision signal. However, since the task-specific output layers differ from the actual high-level layers in the target network, there is bias when optimizing previous sub-nets. To be specific, due to the complex interconnections between different units, the importance of the same unit in previous sub-nets may greatly change with the evolution of the following sub-nets.
To mitigate the bias, we perform progressive re-evaluation of units in preceeding sub-nets. On one hand, we continue to update the discretization functions for preceeding sub-nets during subsequent stages to further prune unimportant units. In practice, we set different pruning intervals, denoted as $e1$ and $e2$, for units in preceding sub-nets and the current sub-net. Notably, $e1$ is set to be larger than $e2$ to primarily focus on pruning in the current sub-net, while allowing longer optimization process before discretization for the previous sub-nets that have already undergone substantial pruning.
@@
$mom(\sigma(\alpha))^{t}=\beta\cdot mom(\sigma(\alpha))^{t-1}+(1-\beta)\cdot\frac{\partial L}{\partial\sigma(\alpha)}^{t-1},$ (6)
where $\beta$ is a smoothing factor and $L$ is the loss. Therefore, the potential of the operation type $T$ in cell $c$, $p_{c}^{T}$, can be calculated as following:
@@ V Experiments
In this section, we evaluate MGAS on CIFAR-10, CIFAR-100 and ImageNet with various baseline methods. We try to answer the following research questions:
@@
    RQ1: Can MGAS discover an architecture with a better accuracy-efficiency trade-off compared with baselines?
@@
    RQ2: Can adaptive pruning consistently achieve a better balance in unit numbers of different granularities for models at different target sizes?
@@
    RQ3: Can multi-stage search save memory without compromising the search quality?
Different variants for each proposed component in MGAS is also discussed within the sub-sections.
@@ V-A Datasets
We evaluate our method on CIFAR-10, CIFAR-100 and ImageNet-1k, which are image classification datasets:
@@
Following the widely adopted ImageNet dataset setting [17, 30], we use only 10% training data during the search to reduce time.
@@ V-B Implementation Details
We follow the configurations of common DARTS-series search spaces [1, 17, 28]. There are two kinds of cells: Normal cells which maintain the resolution and channel number; Reduction cells which reduce the resolution by half and double the channel number. The reduction cells are set at the 1/3 and 2/3 of the depth of the architecture. The candidate operation set consists of 7 operations. And we remove most of the manual constraints on the search space as in [27]: Each edge can preserve an arbitrary number of operations, each node can preserve an arbitrary number of predecessors, and all the cells can have different structures. Moreover, we search for not only the operation types, but also the filter numbers and sparsity patterns in convolution operations.
@@
During the search, we decompose the super-net into three sub-nets according to the feature map resolution, i.e., $n$ is 3, where the later two sub-nets start with the two reduction cells. In each sub-net stage, we prune the network until its parameter number is reduced by half. The pruning intervals for preceding and current sub-nets, $e1$ and $e2$, are set 6 and 2 respectively. The regrowing ratio $r$ is set 0.2. In the fine-tuning process, we prune the network until reaching the target model size. We also apply the Cutout [31] and AutoAugment [32] techniques, as well as the warm-up technique to train the model weights for 5 epochs before the joint optimization with granularity parameters [17, 28]. The search process is conducted with a single NVIDIA A40 GPU. It takes around 2 days to search on CIFAR-10.
@@ V-C Baselines
For fair comparison, we choose baseline methods that are also performed in DARTS-series search spaces.
Most of existing methods fall under the category of coarse-grained methods. Hence, we select eight representative coarse-grained methods with different memory reduction strategies from the past five years: (1) DARTS [1], (2) P-DARTS [28], (3) PC-DARTS [17], (4) GDAS [20], (5) GOLD-NAS [27], (6) MSG-DAS [21], (7) CDARTS [15], (8) MR-DARTS [16], along with one latest fine-grained method: (9) DASS [13] as baselines for CIFAR-10 and CIFAR-100 dataset, as shown in Table I. Among these methods, (1)-(3), (7)-(8) and (9) employ a proxy strategy to save memory, while (4) and (6) apply a decomposition strategy, and (5) directly reduces the candidate operation types.
We further select the baselines which report their model accuracy and model size on ImageNet for comparison on ImageNet dataset, as presented in Table II.
TABLE I: Comparison with SOTA methods on CIFAR-10 and CIFAR-100.
Method Top-1 Accuracy (%) # Params (M) Top-1 Accuracy Density (%/M-params) Granularity
@@
MSG-DAS [21] 97.43 - 3.58±0.09 27.22 - operation
CDARTS [15] 97.52 ± 0.04 84.31 3.9±0.08 25.01 21.62 operation
MR-DARTS [16] 97.51 83.56 4.3 22.68 19.43 operation
MGAS (ours, large) 97.34 83.58 2.1 46.37 39.8 operation, filter, weight
DASS [13] (small) 92.30 67.83 0.7 131.86 96.9 operation, weight
MGAS (ours, small) 96.51 80.61 0.68 141.93 118.54 operation, filter, weight
TABLE II: Comparison with SOTA methods on ImageNet. † indicates directly searching on ImageNet.
Method Accuracy (%) # Params (M) Top-1 Accuracy Density (%/M-params) Granularity
@@
MGAS†(ours, large) 75.5 92.6 4.6 16.41 operation, filter, weight
DASS [13] (small) 48.03 72.72 0.7 68.61 operation, weight
MGAS (ours, small) 65.48 86 0.68 96.29 operation, filter, weight
TABLE III: Memory consumption of MGAS in single-stage and multi-stage settings.
Method Maximum Memory Consumption (G)
Single-stage search 30.7
Stage 1 15.1
Multi-stage search Stage 2 16.1
Stage 3 16.5
Fine-tune 13.3
TABLE IV: Comparison of MGAS with different regrowing ratios.
Regrowing Ratio $r$ Top-1 Accuracy (%) # Params (M) Top-1 Accuracy Density (%/M-params)
0 97.20 2.2 44.18
0.1 97.26 2.2 44.21
0.2 97.34 2.1 46.37
TABLE V: Comparison with SOTA methods in different search spaces on CIFAR10.
@@
Top-1 Accuracy (%) # Params (M) Top-1 Accuracy (%) # Params (M) Top-1 Accuracy (%) # Params (M)
DARTSV1 [1] 96.85 1.23 96.69 1.37 - - Depth-level Proxy
PC-DARTS [17] 97.06 1.68 96.66 1.47 96.56 1.47 Width-level Proxy
GDAS [20] 96.78 0.95 95.34 0.65 96.11 1.13 Decomposition
MGAS 97.04 1.34 97.02 1.25 97.10 1.19 Multi-stage Search
@@
V-D MGAS Discovers Architectures with a Better Accuracy-Efficiency Trade-Off (RQ1)
@@
In this section, we verify whether MGAS can discover architectures with a better accuracy-efficiency trade-off compared with baselines.
V-D1 Results on CIFAR-10 and CIFAR-100
@@
We impletement MGAS with two different search space settings. In order to compare with coarse-grained methods that prioritize accuracy, we start with a search space that contains relatively large models consisting of 14 cells and an initial filter number of 44. On the other hand, to compare with fine-grained methods that prioritize model efficiency and require higher search cost, we explore smaller models with 8 cells and an initial filter number of 24. For fair comparison, we use the same depth and initial width setting for the baseline DASS. We retrain the discovered architecture for 600 epochs for evaluation, and the results are shown in Table I. We utilize top-1 accuracy density [33] to measure the accuracy-efficiency trade-off, which quantifies how efficiently each model uses its parameters by dividing top-1 accuracy by parameter number.
@@
Compared with coarse-grained methods, MGAS achieves comparable accuracy of around 97.4% and 83.5% on CIFAR-10 and CIFAR-100, respectively, using only 2.1M parameters. In the case of CIFAR-10, our method enables a reduction of up to 40% in model size compared to PC-DARTS and MSG-DAS while demonstrating a similar accuracy level with a marginal difference of 0.1%. For CIFAR-100, our method leads to a reduction of up to 45% in model size compared to P-DARTS and MR-DARTS at the same accuracy level. Additionally, MGAS outperforms all the baselines in terms of the top-1 accuracy density, showcasing the advantages of exploring a multi-granularity search space.
In comparison with fine-grained methods, MGAS converges to an architecture with a similar parameter count of approximately 0.7M, yet exhibits substantially higher accuracy. This demonstrates the benefits of balancing the remaining ratios of units at different granularity levels.
It is worth mentioning that the original DASS paper reports a promising result of 95.31% accuracy with only 0.1M parameters on CIFAR-10. However, their search space configuration, which contains models with 20 cells and an initial filter number of 36, differs from ours. As a result, the relative numbers of the remaining operations, filters and weights between their reported result and our implementation are different using the same discretization rules. This discrepancy has a surprisingly significant impact on model performance, underscoring the potential influence of relative unit numbers at different granularity levels. It indicates that methods relying on manually defined remaining ratios cannot guarantee performance without prior knowledge of the search space design, which further emphasizes the necessity of our method.
V-D2 Results on ImageNet

### v4

@@ IV-C Multi-Stage Search
Let the search space of the whole architecture be N. To realize the multi-stage search, we decompose $N$ into sequential n moderate-size search spaces $N_{1},...,N_{i},...,N_{n}$ along the network depth, where $N_{i+1}$ is originally connected with $N_{i}$. In this way, the super-net constructed from the search space is naturally separated into n sub-nets with much smaller sizes. We first coarsely search for the sub-net structures in various stages, removing most of the unimportant units, and then fine-tune the remaining architecture until the target model size is achieved.
@@
Input: dataset $D$, target model size $Param_{min}$, sub-net number $n$, target sub-net size $Param_{min}^{s}$ (s=0,1,,ns=0,1,...,n), regrowing ratio $r$, units in the search space $U$
@@
To supervise the optimization of the sub-net, we add task-specific output layers with predefined structure on top of each sub-net to generate results, so that the ground truth can be directly used as the supervision signal. However, since the task-specific output layers differ from the actual high-level layers in the target network, there is bias when optimizing previous sub-nets. Specifically, due to the complex interconnections between different units, the importance of the same unit in previous sub-nets may greatly change with the evolution of the following sub-nets.
To mitigate the bias, we perform progressive re-evaluation of units in preceeding sub-nets. On one hand, we continue to update the discretization functions for preceeding sub-nets during subsequent stages to further prune unimportant units.
@@
$mom(\sigma(\alpha))^{t}=\gamma\cdot mom(\sigma(\alpha))^{t-1}+(1-\gamma)\cdot\frac{\partial L}{\partial\sigma(\alpha)}^{t-1},$ (6)
where $\gamma$ is a smoothing factor and $L$ is the loss. Therefore, the potential of the operation type $T$ in cell $c$, $p_{c}^{T}$, can be calculated as following:
@@ V Experiments
In this section, we evaluate MG-DARTS on CIFAR-10, CIFAR-100 and ImageNet datasets, comparing it against various baseline methods in both DARTS-like and MobileNet-like search spaces. We try to answer the following research questions:
@@
    RQ1: Can MG-DARTS discover an architecture with a better accuracy-efficiency trade-off compared with baselines?
@@
    RQ2: Can adaptive pruning consistently produce more appropriate retention ratios across different granularity levels for models with varying target sizes?
@@
    RQ3: Can multi-stage search effectively reduce memory consumption without degrading search quality?
Different variants for each proposed component in MG-DARTS is also discussed within the sub-sections.
@@ V-A Datasets
We evaluate our method on CIFAR-10, CIFAR-100 and ImageNet-1k, which are commonly used image classification datasets:
@@
Following the widely adopted ImageNet dataset setting [30], we use only 10% training data during the search to reduce time.
@@ V-B Implementation Details
We evaluate our method on the popular DARTS-like and MobileNet-like search space.
For DARTS search space, we follow the configurations of common settings [1, 30, 41]. There are two kinds of cells: Normal cells which maintain the resolution and channel number; Reduction cells which reduce the resolution by half and double the channel number. The reduction cells are set at the 1/3 and 2/3 of the depth of the architecture. The candidate operation set consists of 7 operations. And we remove most of the manual constraints on the search space as in [40]: Each edge can preserve an arbitrary number of operations, each node can preserve an arbitrary number of predecessors, and all the cells can have different structures. Moreover, we search for not only the operations, but also the filter numbers and weight sparsity patterns in convolution operations.
For MobileNet-like search space, we follow the configurations of SOTA works from the past five years [45]. Since our method is primarily based on a typical differentiable architecture search pipeline, we currently cannot directly search for model depth. Therefore, we fix the depth at 17 layers, following the architecture discovered by WeakNAS [45]. The search space consists of three operations: inverted residual blocks with kernel sizes of $3\times 3$, $5\times 5$, and $7\times 7$. Instead of searching for the rigid expansion ratio, we directly optimize the filter numbers. Additionally, we explore weight sparsity patterns.
@@
During the search, we decompose the super-net into three sub-nets according to the feature map resolution, i.e., $n$ is 3, where the later two sub-nets start with the two reduction cells. The regrowing ratio $r$ is set 0.2. In the architecture fine-tuning stage, we prune the network until reaching the target model size. To accelerate the search process at each stage, we initialize the thresholds $t_{g}$ using the minimum value of the current granularity parameters $g$, rather than starting from zero. We also apply the Cutout [46] and AutoAugment [47] techniques, as well as the warm-up process of training the model weights for 5 epochs before the joint optimization with granularity parameters [30, 41]. The search process is conducted with a single NVIDIA A40 GPU. It takes around 1 day to search on both CIFAR-10 in DARTS search space and ImageNet in MobileNet-like search space.
@@ V-C Baselines
For the DARTS-like search space, most of existing methods are differentiable and only conduct operation-level search. Hence, we select 15 representative single-granularity methods [1, 41, 30, 20, 40, 48, 25, 49, 50, 37, 22, 51, 36, 35, 33], along with 2 latest multi-granularity methods [17, 14] as baselines for CIFAR-10 and CIFAR-100 datasets, as shown in Table I. Among them, since DASS [17] is the most similar one to our method and is originally designed for extremely compact networks, we generate a small-size architecture using our approach and re-implement DASS under our customized search configuration to ensure a fair comparison.
For the ImageNet dataset, we include only those baselines that report both model accuracy and model size, as summarized in Table II.
For the MobileNet-like search space, most of existing methods are non-differentiable and conduct operation-level and filter-level search. We compare with several state-of-the-art methods [19, 48, 45, 51], as presented in Table III.
TABLE I: Comparison with SOTA methods on CIFAR-10 and CIFAR-100 in DARTS-like search space.
Method Top-1 Accuracy (%) Params (M) Accuracy Density (%/M-params) Granularity
@@
CTNAS [48] 97.41 - 3.6 27.05 - operation
CDARTS [25] 97.52 84.31 3.9 25.01 21.62 operation
Shapley-NAS [49] 97.53 - 3.6 27.09 - operation
$\beta$-DARTS [50] 97.47 83.76 3.75 25.99 22.33 operation
PA-DARTS [37] 97.52 83.46 3.57 27.31 23.71 operation
U-DARTS [22] 97.41 83.68 3.3 29.51 25.35 operation
Zero-Cost-PT [51] 97.38 - 4.6 21.16 - operation
$l$-DARTS [36] 97.54 83.62 4.1 23.79 20.39 operation
ACA-PC-DARTS [35] 97.46 - 3.6 27.07 - operation
CSC-DARTS [33] 97.48 82.18 4.6 21.19 17.86 operation
TangleNAS [14] 97.44 - 3.58 27.21 - operation, filter
MG-DARTS (ours, large) 97.47 83.79 2.2 44.3 38.08 operation, filter, weight
DASS [17] (small) 95.49 77.09 0.71 133.92 108.57 operation, weight
MG-DARTS (ours, small) 96.84 81.76 0.7 138.34 116.8 operation, filter, weight
TABLE II: Comparison with SOTA methods on ImageNet in DARTS-like search space.
Method Accuracy (%) Params (M) Accuracy Density (%/M-params) Granularity
@@
Shapley-NAS [49] 76.1 - 5.4 14.09 operation
$\beta$-DARTS [50] 76.1 93 5.5 13.83 operation
PA-DARTS [37] 75.3 92.25 5.2 14.48 operation
U-DARTS [22] 73.9 91.9 4.9 15.08 operation
Zero-Cost-PT [51] 75.4 92.4 6.3 11.96 operation
$l$-DARTS [36] 75.09 92.35 5.8 12.94 operation
CSC-DARTS [33] 74.5 91.7 4.6 16.19 operation
TangleNAS [14] 74.3 - 5.15 14.42 operation, filter
MG-DARTS (ours, large) 75.3 92 4 18.82 operation, filter, weight
DASS [17] (small) 63.79 84.74 0.71 89.84 operation, weight
MG-DARTS (ours, small) 64.89 86.36 0.7 92.7 operation, filter, weight
TABLE III: Comparison with SOTA methods on ImageNet in MobileNet-like search space.
Method Top-1 Accuracy (%) Params (M) Accuracy Density (%/M) Granularity
CTNAS [48] 77.3 operation, filter
WeakNAS [45] 81.3 8.8 8.77 operation, filter
Zero-Cost-PT [51] 76.4 8.0 9.55 operation, filter
TangleNAS [14] 77.4 7.6 10.18 operation, filter
MG-DARTS 77.2 7.4 10.43 operation, filter, weight
TABLE IV: Comparison of MG-DARTS in single-stage and multi-stage settings.
Method Maximum Memory Consumption (G) Search Cost (h) Accuracy (%)
Single-stage 26.4 8 96.56
Stage 1 12.1 10 96.56
Multi-stage Stage 2 12.4
Stage 3 11.9
Fine-tune 7.8
TABLE V: Comparison of MG-DARTS with different regrowing ratios.
Regrowing Ratio $r$ Top-1 Accuracy (%) Params (M) Accuracy Density (%/M-params)
0 97.14 2.4 40.47
0.1 97.21 2.4 40.5
0.2 97.47 2.2 44.30
TABLE VI: Comparison with SOTA methods in different DARTS-like search spaces on CIFAR10.
@@
Top-1 Accuracy (%) Params (M) Top-1 Accuracy (%) Params (M) Top-1 Accuracy (%) Params (M)
DARTSV1 [1] 96.85 1.23 96.69 1.37 - - Depth-level
PC-DARTS [30] 97.06 1.68 96.66 1.47 96.56 1.47 Width-level
GDAS [20] 96.78 0.95 95.34 0.65 96.11 1.13 Path-level
MG-DARTS 97.04 1.34 97.02 1.25 97.10 1.19 Depth-level with Re-evaluation
@@
V-D MG-DARTS Discovers Architectures with a Better Accuracy-Efficiency Trade-Off (RQ1)
@@
In this section, we verify whether MG-DARTS can discover architectures with a better accuracy-efficiency trade-off compared with baselines.
V-D1 Results on CIFAR-10 and CIFAR-100 in DARTS search space
@@
We implement MG-DARTS with two different search space settings. In order to compare with most of single-granularity methods that prioritize accuracy, we adopt a large search space configuration, which includes relatively large models with 14 cells and an initial filter number of 44. On the other hand, to compare with most of multi-granularity methods that prioritize model efficiency and require higher search cost, we adopt a small search space configuration, exploring more compact models with 8 cells and an initial filter number of 24. For fair comparison, we adopt the same depth and initial width settings for the baseline DASS. We set the weight retention ratio to 90% for DASS during pruning, as this ratio shows the most promising results under our search configuration according to our toy ablation study (Table VII). The results are shown in Table I. We utilize accuracy density [52] to measure the accuracy-efficiency trade-off, which quantifies how efficiently each model uses its parameters by dividing top-1 accuracy by parameter number.
@@
In the large model setting, MG-DARTS achieves comparable accuracy of around 97.5% and 83.8% on CIFAR-10 and CIFAR-100, respectively, using only 2.2M parameters. On CIFAR-10, MG-DARTS reduces model size by an average of up to 40% compared to baselines with a marginal accuracy difference of 0.05%. For CIFAR-100, MG-DARTS outperforms all baselines in accuracy except CDARTS, while achieving up to a 43% reduction in model size compared to CDARTS, with only a modest accuracy gap of around 0.5%. Additionally, MG-DARTS outperforms all the baselines in terms of the accuracy density, showcasing the advantages of exploring a comprehensive multi-granularity search space.
In the small model setting, MG-DARTS exhibits a higher accuracy than DASS while maintaining a similar parameter count around 0.7M.
It is worth mentioning that in the original DASS work, an impressive 95.31% accuracy is achieved on CIFAR-10 with only 0.1M parameters. However, under our 0.7M parameter setting, DASS only reaches 95.49% accuracy, suggesting that it may not be fully leveraging its realizing under our experimental configuration.
We attribute this discrepancy to two main factors. First, we adopt the same pretraining schedule as reported in the original paper, i.e. 50 epochs of super-net training and 20 epochs of pruning, which was designed for a smaller 16-channel super-net. In our 24-channel setting, the super-net likely requires longer training to converge, potentially leading to suboptimal pruning masks. Second, to ensure a fair comparison with our method, we retrain all discovered architectures from scratch, whereas original DASS inherits pretrained weights to restore accuracy during evaluation.
TABLE VII: Comparison of DASS and MG-DARTS under different search and evaluation settings. MG-DARTS is less sensitive to the degree of super-net optimization.
Method Search Epoch Evaluation Mode Weight Ratio (%) Accuracy (%) Search Cost (h)
DASS 10 from scratch 30 73.44 7
50 80.36
70 84.68
90 86.58
96.3 86.82
100 86.35
finetune 50 79.62
90 86.43
96.3 86.52
100 86.75
70 from scratch 50 86.30 55
90 87.01
100 86.66
finetune 50 87.12
90 87.78
100 87.57
MG-DARTS 10 from scratch Adaptive (96.3) 93.37 5
To support this analysis, we conduct controlled experiments on CIFAR-10 using DASS under various search and evaluation settings. We vary the weight retention ratio (30%, 50%, 70%, and 90%) while adjusting the number of operations to keep the overall model size approximately constant around 0.7M. Additionally, we compare different numbers of search epochs (10 and 70) to simulate varying levels of super-net optimization. For evaluation, we consider both training from scratch and fine-tuning based on the pretrained super-net weights.
As shown in Table VII, DASS is highly sensitive to the retention ratio when the super-net is insufficiently optimized. With only 10 search epochs, models with lower retention ratios (e.g., 30% and 50%) suffer significant performance drops, likely due to suboptimal sparsity patterns. Fine-tuning based on pretrained weights during evaluation partially mitigates this issue in sparse regimes (e.g., a 6.18% improvement at 50% ratio). Increasing the search epochs to 70 also alleviates this problem, particularly under high sparsity levels (e.g., a 5.94% gain at 50% ratio). These results confirm DASS’s reliance on pretraining.
In contrast, MG-DARTS achieves promising result when trained from scratch even without extensive pretraining. To eliminate the potential influence of search space topology constraints for fair comparison, we also apply the adaptive retention ratio obtained by MG-DARTS to DASS. This yields the best scratch-trained result at 86.82% among all DASS settings, and performs only slightly worse than the fully dense model when fine-tuned. These results suggest that MG-DARTS, by adaptively determining retention ratios at multiple granularity levels, is more robust to under-optimized super-nets and inaccurate architectural patterns.
Overall, the performance gap reflects two fundamentally different design philosophies: DASS emphasizes aggressive sparsity through heavy pretraining, whereas MG-DARTS promotes resilience by adaptively controlling retention across multiple granularity levels.
V-D2 Results on ImageNet in DARTS search space